OpenStudy (anonymous):

Simplify this expression g-3(f-g)-2f

OpenStudy (nikato):

distribute the -3(f-g)

OpenStudy (nikato):

ditribute is like this a(b-c)=ab-ac 0r a(b+c)=ab+ac